If an engine does 300J of net work and exhausts 800J of heat per cycle, what is its thermal efficiency? ?=   %
Answer #1

Thermal Efficiency (in %) =\frac{Work Done}{Total Energy Exhausted}*100=\frac{300}{800+300}*100=27.27%